# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 6.3. Characterizing small type
The aim of this section is to explicitly describe the non-selfdual Wadge classes of small type (that is, type $1$ or $2$), at least in the case of countable level (see Theorem 6.11). We will often consider $\mathbf{\Delta}=\bigcup_{n\in\omega}(\mathbf{\Gamma}_{n}\cup\widecheck{\mathbf{\Gamma}}_{n})$, where the $\mathbf{\Gamma}_{n}$ are intended to be the predecessors of a certain class $\mathbf{\Gamma}$. We remark that this is simply a convenient way to unify the cases in which $\mathbf{\Gamma}$ is a successor (all $\mathbf{\Gamma}_{n}$ are the same) and a countable limit ($\mathbf{\Gamma}_{0}\subsetneq\mathbf{\Gamma}_{1}\subsetneq\cdots$).
@@
The case $\xi=0$ can be easily obtained as a consequence of Lemma 6.9. The general case will then follow, using the methods of Section 3.4 in conjunction with Corollary 6.4 and Lemma 6.10. ∎
Chapter 7 Wadge theory: closure and preservation
This chapter is the Wadge-theoretic core of this monograph. In fact, the results of Section 7.5 show that suitable assumptions concerning the level, type and separation property will ensure closure under separated unions. These closure properties will be an essential ingredient for our existence results (see Theorems 9.13 and 10.3). Along the way, we will show that the operation of separated differences preserves the property of having small type and the separation property. The results concerning good Wadge classes (which were originally obtained in [CMM1]) will ultimately ensure that Steel’s Theorem 8.9 can be applied to the Wadge classes that are relevant to our classification (see Theorem 9.7 and Lemma 9.8).
@@ 7.1. Closure properties: preliminaries
As will become increasingly clear, what we need from Wadge theory are closure properties. These properties are spelled out in the following definition, while the subsequent lemmas will be useful later on, towards obtaining the more sophisticated closure properties of Section 7.5.
@@
The next four results show that these closure properties are well-behaved with respect to relativization and expansion.
@@
by Lemmas 3.29 and 3.30, as desired. ∎
We conclude this section with two elementary results in the same vein (the straightforward proof of Lemma 7.6 is left to the reader).
@@ 7.2. Closure properties: good Wadge classes
The following key notion (see [CMM1, Definition 12.1]) is essentially due to van Engelen, although he did not give it a name.111 One important difference is that van Engelen’s treatment of this notion is fundamentally tied to Louveau’s classification of the Borel Wadge classes from [Lo1], hence it is limited to the Borel context. Good Wadge classes will enable us to define a well-behaved notion of exact topological complexity (see Section 9.1). Furthermore, they will ultimately allow us to apply Theorem 8.9 (since homogeneous spaces have good complexity by Theorem 9.7, and good Wadge classes are reasonably closed by Lemma 9.8).
@@
First observe that $\mathbf{\Gamma}(Z)$ and $\mathbf{\Gamma}(W)$ are non-selfdual by Theorem 3.13. Furthermore, Theorem 3.20 guarantees that $\ell(\mathbf{\Gamma}(Z))\geq 1$ iff $\ell(\mathbf{\Gamma}(W))\geq 1$. Therefore, by Proposition 7.9, it remains to show that $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(Z))\subseteq\mathbf{\Gamma}(Z)$ for every $n\in\omega$ iff $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(W))\subseteq\mathbf{\Gamma}(W)$ for every $n\in\omega$. In order to see this, use the methods of [CMM2, Section 9] and [CMM2, Lemma 21.5] to obtain $\mathbf{\Lambda}_{n}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$ for $n\in\omega$ such that each $\mathbf{\Lambda}_{n}(Z)=\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(Z))$ and each $\mathbf{\Lambda}_{n}(W)=\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(W))$. Then, the desired equivalence will follow from Theorem 3.14. ∎
The remainder of this section is devoted to obtaining the fundamental closure properties of good Wadge classes (see Corollary 7.13), as a particular case of a result of independent interest (see Theorem 7.12). We remark that, for the uniqueness part of the classification, no further Wadge-theoretic analysis is required. However, in order to obtain the existence results of Sections 9.4 and 10.2, we will need the more sophisticated closure properties discussed in Section 7.5.
The following result first appeared as [CMM1, Lemma 12.3], and it generalizes [AHN, Lemma 3.6.a]. However, we will give a new proof, which relies on separated differences and Theorem 3.35 instead of [AHN]. Corollary 7.13 first appeared as [CMM1, Theorem 12.4].
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, and let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$. Assume that $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{1}(Z))\subseteq\mathbf{\Gamma}$ whenever $1\leq n<\omega$. Then $\mathbf{\Gamma}$ has the following closure properties:
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, let $\xi<\omega_{1}$, and let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$. Assume that $\ell(\mathbf{\Gamma})\geq\xi$ and $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{1+\xi}(Z))\subseteq\mathbf{\Gamma}$ whenever $1\leq n<\omega$. Then $\mathbf{\Gamma}$ has the following closure properties:
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, and let $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(Z)$ be a good Wadge class in $Z$. Then $\mathbf{\Gamma}$ has the following closure properties:
@@ 7.3. Preservation of small type under separated differences
The aim of this section is to show that the operation $\mathsf{SD}_{\eta}(\mathbf{\Delta},-)$ preserves the property of having small type for classes of level 00 (see Theorem 7.15 for a precise statement). This is one of the ingredients needed to obtain the crucial closure properties of Section 7.5. We begin with a technical lemma.
@@
Since $\mathbf{\Gamma}_{n}\subseteq\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast}))$ for each $n$, it is clear that $\mathbf{\Gamma}_{n}^{\ast}\subseteq\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast})$ for each $n$. Similarly, using Lemma 3.33 and the assumption that $\mathbf{\Delta}$ is selfdual, one sees that $\widecheck{\mathbf{\Gamma}}_{n}\subseteq\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast}))$ for each $n$. Since $\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast}))$ is closed under $\mathsf{SU}_{0}$ by Lemmas 7.6 and 7.7, the inclusion $\subseteq$ follows.
@@
Set $V_{-1}=\bigcup\{U_{\mu,n}:\mu<\eta\text{ and }n\in\omega\}$. By Theorem 2.4, there exist $V^{\prime}_{k}\in\mathbf{\Sigma}^{0}_{1}(Z)$ for $-1\leq k<\omega$ satisfying the following conditions:
  • $V^{\prime}_{k}\subseteq V_{k}$ for $-1\leq k<\omega$,
@@
$\mathbf{\Gamma}^{\ast}=\mathsf{SU}_{0}\left(\bigcup_{n\in\omega}(\mathbf{\Gamma}^{\ast}_{n}\cup\widecheck{\mathbf{\Gamma}}^{\ast}_{n})\right)\text{ or }\widecheck{\mathbf{\Gamma}}^{\ast}=\mathsf{SU}_{0}\left(\bigcup_{n\in\omega}(\mathbf{\Gamma}^{\ast}_{n}\cup\widecheck{\mathbf{\Gamma}}^{\ast}_{n})\right).$
@@
$\mathbf{\Gamma}=\mathsf{SU}_{0}\left(\bigcup_{n\in\omega}(\mathbf{\Gamma}_{n}\cup\widecheck{\mathbf{\Gamma}}_{n})\right)\text{ or }\widecheck{\mathbf{\Gamma}}=\mathsf{SU}_{0}\left(\bigcup_{n\in\omega}(\mathbf{\Gamma}_{n}\cup\widecheck{\mathbf{\Gamma}}_{n})\right)$
by Lemmas 3.33 and 7.14, the desired conclusion follows from Theorem 6.11. ∎
@@ 7.4. Preservation of the separation property under separated differences
The aim of this section is to show that the operation $\mathsf{SD}_{\eta}(\mathbf{\Delta},-)$ preserves the separation property (see Theorem 7.18 for a precise statement). This is one of the ingredients needed to obtain the crucial closure properties of Section 7.5. We begin with a technical lemma.
@@
[…]
@@
However, the case $\mathbf{\Gamma}^{\ast}=\{\varnothing\}$ contradicts the assumption that $\mathbf{\Gamma}^{\ast}$ has the separation property. Furthermore, since each $\widecheck{\mathsf{D}}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$ has the separation property by Lemmas 3.34, 6.8 and 7.16, the case $\mathbf{\Gamma}^{\ast}=\mathsf{D}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$ would lead to a contradiction by Theorem 5.8. By Lemma 6.8, in order to conclude the proof, it will be enough to show that $\mathbf{\Gamma}=\widecheck{\mathsf{SU}}_{0}(\mathbf{\Delta}^{\prime})$ for some selfdual $\mathbf{\Delta}^{\prime}\subseteq\mathcal{P}(Z)$. In the case $\mathbf{\Gamma}^{\ast}=\{Z\}$, this clearly follows from Lemmas 3.33 and 7.16. Now assume that $\mathbf{\Gamma}^{\ast}=\widecheck{\mathsf{D}}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$, where $1\leq m<\omega$. Observe that $t(\mathbf{\Gamma}^{\ast})=1$ by Theorem 3.8. Furthermore, it follows from Lemma 3.34 and Theorem 3.35 that $\ell(\mathbf{\Gamma}^{\ast})=0$. By Theorems 6.11 and 7.15, this means that there exists a selfdual $\mathbf{\Delta}^{\prime}\subseteq\mathcal{P}(Z)$ such that either $\mathbf{\Gamma}=\mathsf{SU}_{0}(\mathbf{\Delta}^{\prime})$ or $\mathbf{\Gamma}=\widecheck{\mathsf{SU}}_{0}(\mathbf{\Delta}^{\prime})$. Assume, in order to get a contradiction, that $\mathbf{\Gamma}=\mathsf{SU}_{0}(\mathbf{\Delta}^{\prime})$. Notice that $\mathsf{D}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$ is closed under $\mathsf{SU}_{0}$ by Lemmas 3.34, 7.6 and 7.16. It follows from Lemmas 3.33 and 7.7 that $\widecheck{\mathbf{\Gamma}}$ is closed under $\mathsf{SU}_{0}$. Hence
@@ 7.5. Closure properties: the main results
This section is the culmination (and the conclusion) of the purely Wadge-theoretic portion of this article.
@@
Using the methods of Section 3.4 and Lemma 7.5, one sees that the case $\xi=0$ implies the general case. So assume, in order to get a contradiction, that the desired result does not hold in this particular case. By Theorem 3.4, we can assume that $\mathbf{\Gamma}$ is a $\subseteq$-minimal counterexample (in other words, $\ell(\mathbf{\Gamma})=0$ and $\mathbf{\Gamma}$ does not have the separation property, but $\mathbf{\Gamma}$ is not closed under $\mathsf{SU}_{0}$). Fix $\eta$, $\mathbf{\Delta}$ and $\mathbf{\Gamma}^{\ast}$ as given by Theorem 3.35 such that $\mathbf{\Gamma}=\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathbf{\Gamma}^{\ast})$. Notice that $\mathbf{\Gamma}^{\ast}$ does not have the separation property, otherwise $\mathbf{\Gamma}$ would have the separation property by Theorem 7.18. Similarly, Lemma 7.7 shows that $\mathbf{\Gamma}^{\ast}$ is not closed under $\mathsf{SU}_{0}$. Furthermore, as $\mathbf{\Delta}$ is selfdual and $\mathbf{\Gamma}^{\ast}\subseteq\mathbf{\Delta}\subseteq\mathbf{\Gamma}$, one sees that $\mathbf{\Gamma}^{\ast}\subsetneq\mathbf{\Gamma}$. It follows that $\ell(\mathbf{\Gamma}^{\ast})\geq 1$, otherwise the minimality of $\mathbf{\Gamma}$ would be contradicted. In particular, $\mathbf{\Gamma}^{\ast}$ is closed under $\mathsf{SU}_{0}$ (notice that $\mathbf{\Gamma}^{\ast}\neq\{Z\}$ because $\mathbf{\Gamma}^{\ast}$ does not have the separation property), which is a contradiction. ∎
@@
Using the methods of Section 3.4, Corollary 6.4 and Lemma 7.5, one sees that the case $\xi=0$ implies the general case. So assume, in order to get a contradiction, that the desired result does not hold in this particular case. By Theorem 3.4, we can assume that $\mathbf{\Gamma}$ is a $\subseteq$-minimal counterexample (in other words, $\ell(\mathbf{\Gamma})=0$ and $t(\mathbf{\Gamma})=3$, but $\mathbf{\Gamma}$ is not closed under $\mathsf{SU}_{0}$). Fix $\eta$, $\mathbf{\Delta}$ and $\mathbf{\Gamma}^{\ast}$ as given by Theorem 3.35 such that $\mathbf{\Gamma}=\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathbf{\Gamma}^{\ast})$. Observe that $t(\mathbf{\Gamma}^{\ast})\neq 0$, otherwise we would have $t(\mathbf{\Gamma})\in\{1,2\}$ by Lemma 3.33, Theorem 6.11 and Lemma 7.17. To conclude the proof, we will show that the cases $\ell(\mathbf{\Gamma}^{\ast})=0$ and $\ell(\mathbf{\Gamma}^{\ast})\geq 1$ both lead to a contradiction.

### v3

@@ 4.3. Characterizing small type
The aim of this subsection is to explicitly describe the non-selfdual Wadge classes of small type (that is, type $1$ or $2$), at least in the case of countable level (see Theorem 4.11). We will often consider $\mathbf{\Delta}=\bigcup_{n\in\omega}(\mathbf{\Gamma}_{n}\cup\widecheck{\mathbf{\Gamma}}_{n})$, where the $\mathbf{\Gamma}_{n}$ are intended to be the predecessors of a certain class $\mathbf{\Gamma}$. We remark that this is simply a convenient way to unify the cases in which $\mathbf{\Gamma}$ is a successor (all $\mathbf{\Gamma}_{n}$ are the same) and a countable limit ($\mathbf{\Gamma}_{0}\subsetneq\mathbf{\Gamma}_{1}\subsetneq\cdots$).
@@
The case $\xi=0$ can be easily obtained as a consequence of Lemma 4.9. The general case will then follow, using the methods of §3.4 in conjunction with Corollary 4.4 and Lemma 4.10. ∎
5. Wadge theory: closure and preservation
@@ 5.1. Closure properties: preliminaries
As will become increasingly clear, what we need from Wadge theory are closure properties. These properties are spelled out in the following definition, while the subsequent lemmas will be useful later on, towards obtaining the more sophisticated closure properties of §5.5.
@@
The next four results show that these closure properties are well-behaved with respect to relativization and expansions.
@@
by Lemmas 3.25 and 3.26, as desired. ∎
We conclude this subsection with two elementary results in the same vein (the straightforward proof of Lemma 5.6 is left to the reader).
@@ 5.2. Closure properties: good Wadge classes
The following key notion (see [CMM1, Definition 12.1]) is essentially due to van Engelen, although he did not give it a name.171717 One important difference is that van Engelen’s treatment of this notion is fundamentally tied to Louveau’s classification of the Borel Wadge classes from [Lo1], hence it is limited to the Borel context. Good Wadge classes will enable us to define a well-behaved notion of exact topological complexity (see §6.1). Furthermore, the closure properties given by Corollary 5.13 will be applied directly in the proofs of Theorems 6.12 and 7.3.
@@
First observe that $\mathbf{\Gamma}(Z)$ and $\mathbf{\Gamma}(W)$ are non-selfdual by Theorem 3.11. Furthermore, Theorem 3.17 guarantees that $\ell(\mathbf{\Gamma}(Z))\geq 1$ iff $\ell(\mathbf{\Gamma}(W))\geq 1$. Therefore, by Proposition 5.9, it remains to show that $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(Z))\subseteq\mathbf{\Gamma}(Z)$ for every $n\in\omega$ iff $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(W))\subseteq\mathbf{\Gamma}(W)$ for every $n\in\omega$. In order to see this, use the methods of [CMM2, §9] and [CMM2, Lemma 21.5] to obtain $\mathbf{\Lambda}_{n}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$ for $n\in\omega$ such that each $\mathbf{\Lambda}_{n}(Z)=\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(Z))$ and each $\mathbf{\Lambda}_{n}(W)=\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{2}(W))$. Then, the desired equivalence will follow from Theorem 3.12. ∎
The remainder of this subsection is devoted to obtaining the fundamental closure properties of good Wadge classes (see Corollary 5.13), as a particular case of a result of independent interest (see Theorem 5.12). The following result first appeared as [CMM1, Lemma 12.3], and it generalizes [AHN, Lemma 3.6.a]. However, we will give a new proof, which relies on separated differences and Theorem 3.55 instead of [AHN]. Corollary 5.13 first appeared as [CMM1, Theorem 12.4].
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, and let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$. Assume that $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{1}(Z))\subseteq\mathbf{\Gamma}$ whenever $1\leq n<\omega$. Then:
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, let $\xi<\omega_{1}$, and let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$. Assume that $\ell(\mathbf{\Gamma})\geq\xi$ and $\mathsf{D}_{n}(\mathbf{\Sigma}^{0}_{1+\xi}(Z))\subseteq\mathbf{\Gamma}$ whenever $1\leq n<\omega$. Then:
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, and let $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(Z)$ be a good Wadge class in $Z$. Then:
@@ 5.3. Preservation of small type under separated differences
The aim of this subsection is to show that the operation $\mathsf{SD}_{\eta}(\mathbf{\Delta},-)$ preserves the property of having small type for classes of level 0 (see Theorem 5.15 for a precise statement). This is one of the ingredients needed to obtain the crucial closure properties of §5.5. We begin with a technical lemma.
@@
Since $\mathbf{\Gamma}_{n}\subseteq\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast}))$ for each $n$, it is clear that $\mathbf{\Gamma}_{n}^{\ast}\subseteq\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast})$ for each $n$. Similarly, using Lemma 3.53 and the assumption that $\mathbf{\Delta}$ is selfdual, one sees that $\widecheck{\mathbf{\Gamma}}_{n}\subseteq\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast}))$ for each $n$. Since $\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathsf{SU}_{0}(\mathbf{\Delta}\!^{\ast}))$ is closed under $\mathsf{SU}_{0}$ by Lemmas 5.6 and 5.7, the inclusion $\subseteq$ follows.
@@
Set $V_{-1}=\bigcup\{U_{\mu,n}:\mu<\eta\text{ and }n\in\omega\}$. By Theorem 2.3, there exist $V^{\prime}_{k}\in\mathbf{\Sigma}^{0}_{1}(Z)$ for $1\leq k<\omega$ satisfying the following conditions:
  • $V^{\prime}_{k}\subseteq V_{k}$ for $1\leq k<\omega$,
@@
[…]
@@
[…]
by Lemmas 3.53 and 5.14, the desired conclusion follows from Theorem 4.11. ∎
@@ 5.4. Preservation of the separation property under separated differences
The aim of this subsection is to show that the operation $\mathsf{SD}_{\eta}(\mathbf{\Delta},-)$ preserves the separation property (see Theorem 5.18 for a precise statement). This is one of the ingredients needed to obtain the crucial closure properties of §5.5.
@@
where the equality holds by Lemma 5.7, since $\{\varnothing\}$ is obviously closed under $\mathsf{SU}_{0}$.
@@
However, the case $\mathbf{\Gamma}^{\ast}=\{\varnothing\}$ contradicts the assumption that $\mathbf{\Gamma}^{\ast}$ has the separation property. Furthermore, since each $\widecheck{\mathsf{D}}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$ has the separation property by Lemmas 3.54, 4.8 and 5.16, the case $\mathbf{\Gamma}^{\ast}=\mathsf{D}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$ would lead to a contradiction by Theorem 3.50. By Lemma 4.8, in order to conclude the proof, it will be enough to show that $\mathbf{\Gamma}=\widecheck{\mathsf{SU}}_{0}(\mathbf{\Delta}^{\prime})$ for some selfdual $\mathbf{\Delta}^{\prime}\subseteq\mathcal{P}(Z)$. In the case $\mathbf{\Gamma}^{\ast}=\{Z\}$, this clearly follows from Lemmas 3.53 and 5.16. Now assume that $\mathbf{\Gamma}^{\ast}=\widecheck{\mathsf{D}}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$, where $1\leq m<\omega$. Observe that $t(\mathbf{\Gamma}^{\ast})=1$ by Theorem 3.6. Furthermore, it follows from Lemma 3.54 and Theorem 3.55 that $\ell(\mathbf{\Gamma}^{\ast})=0$. By Theorems 4.11 and 5.15, this means that there exists a selfdual $\mathbf{\Delta}^{\prime}\subseteq\mathcal{P}(Z)$ such that either $\mathbf{\Gamma}=\mathsf{SU}_{0}(\mathbf{\Delta}^{\prime})$ or $\mathbf{\Gamma}=\widecheck{\mathsf{SU}}_{0}(\mathbf{\Delta}^{\prime})$. Assume, in order to get a contradiction, that $\mathbf{\Gamma}=\mathsf{SU}_{0}(\mathbf{\Delta}^{\prime})$. Notice that $\mathsf{D}_{m}(\mathbf{\Sigma}^{0}_{1}(Z))$ is closed under $\mathsf{SU}_{0}$ by Lemmas 3.54, 5.6 and 5.16. It follows from Lemmas 3.53 and 5.7 that $\widecheck{\mathbf{\Gamma}}$ is closed under $\mathsf{SU}_{0}$. Hence
@@ 5.5. Closure properties: the main results
This subsection is the culmination (and the conclusion) of the purely Wadge-theoretic portion of this article.
@@
Using the methods of §3.4 and Lemma 5.5, one sees that the case $\xi=0$ implies the general case. So assume, in order to get a contradiction, that the desired result does not hold in this particular case. By Theorem 3.4, we can assume that $\mathbf{\Gamma}$ is a $\subseteq$-minimal counterexample (in other words, $\ell(\mathbf{\Gamma})=0$ and $\mathbf{\Gamma}$ does not have the separation property, but $\mathbf{\Gamma}$ is not closed under $\mathsf{SU}_{0}$). Fix $\eta$, $\mathbf{\Delta}$ and $\mathbf{\Gamma}^{\ast}$ as given by Theorem 3.55 such that $\mathbf{\Gamma}=\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathbf{\Gamma}^{\ast})$. Notice that $\mathbf{\Gamma}^{\ast}$ does not have the separation property, otherwise $\mathbf{\Gamma}$ would have the separation property by Theorem 5.18. Similarly, Lemma 5.7 shows that $\mathbf{\Gamma}^{\ast}$ is not closed under $\mathsf{SU}_{0}$. Furthermore, as $\mathbf{\Delta}$ is selfdual and $\mathbf{\Gamma}^{\ast}\subseteq\mathbf{\Delta}\subseteq\mathbf{\Gamma}$, one sees that $\mathbf{\Gamma}^{\ast}\subsetneq\mathbf{\Gamma}$. It follows that $\ell(\mathbf{\Gamma}^{\ast})\geq 1$, otherwise the minimality of $\mathbf{\Gamma}$ would be contradicted. In particular, $\mathbf{\Gamma}^{\ast}$ is closed under $\mathsf{SU}_{0}$ (notice that $\mathbf{\Gamma}^{\ast}\neq\{Z\}$ because $\mathbf{\Gamma}^{\ast}$ does not have the separation property), which is a contradiction. ∎
@@
Using the methods of §3.4, Corollary 4.4 and Lemma 5.5, one sees that the case $\xi=0$ implies the general case. So assume, in order to get a contradiction, that the desired result does not hold in this particular case. By Theorem 3.4, we can assume that $\mathbf{\Gamma}$ is a $\subseteq$-minimal counterexample (in other words, $\ell(\mathbf{\Gamma})=0$ and $t(\mathbf{\Gamma})=3$, but $\mathbf{\Gamma}$ is not closed under $\mathsf{SU}_{0}$). Fix $\eta$, $\mathbf{\Delta}$ and $\mathbf{\Gamma}^{\ast}$ as given by Theorem 3.55 such that $\mathbf{\Gamma}=\mathsf{SD}_{\eta}(\mathbf{\Delta},\mathbf{\Gamma}^{\ast})$. Observe that $t(\mathbf{\Gamma}^{\ast})\neq 0$, otherwise we would have $t(\mathbf{\Gamma})\in\{1,2\}$ by Lemma 3.53, Theorem 4.11 and Lemma 5.17. To conclude the proof, we will show that the cases $\ell(\mathbf{\Gamma}^{\ast})=0$ and $\ell(\mathbf{\Gamma}^{\ast})\geq 1$ both lead to a contradiction.